OpenStudy (anonymous):

Find all the real zeros of the polynomial P(x)=x^4 - 3 x^3 - 6 x^2 + 8 x .

OpenStudy (whpalmer4):

Can you factor the polynomial? Once you've factored it into the form \[P(x) = (x-a)(x-b)(x-c)(x-d)\]a-d are the roots.

OpenStudy (anonymous):

thank you

OpenStudy (whpalmer4):

Post your answers and I'll check them for you if you want.

OpenStudy (anonymous):

this is as far as i was able to get x(x^2-3x^2-6x+8) can someone please help me?

OpenStudy (anonymous):

have you tried (x-1) as a root?

OpenStudy (anonymous):

no because it was all tied up in a poly and I thought i had to factor first to break it down

OpenStudy (anonymous):

Divide \(x(x^3-3x^2-6x+8)\) with \((x-1)\) then tell me what you got :)

OpenStudy (anonymous):

I don't even know how to compute that.

OpenStudy (anonymous):

Use synthetic division; have you learned that yet?

OpenStudy (anonymous):

yes. i'm just confuse with that x in the front.

OpenStudy (anonymous):

you can ignore that when you do the synthetic division, and then multiply it back afterwards. just divide \((x^3−3x^2−6^x+8)\) with \((x-1)\)

OpenStudy (anonymous):

so when when I did the division my numbers were 1, 4, 10, 2 there wasn't a zero in there. I tried it with the other factors of 8 and 4 seemed to work it gave me 1, 1, -2, 0 but i'm not sure what to do with that now

OpenStudy (anonymous):

it's ok.when you divide \(x^3-3x^2-6^x+8\) with \(x-1\) you should get \(x^2-2x-8\), so the whold equation becomes \[x(x-1)(x^2-2x-8)=0\]right?

OpenStudy (anonymous):

I didn't get those numbers but i probably did it wrong

OpenStudy (anonymous):

okay, yes i just did

OpenStudy (anonymous):

do i need to take the 1 and create another division with the new numbers of 1, -2, 8 and 0 now?

OpenStudy (anonymous):

well no zero bc there wasn't one when i did it

OpenStudy (anonymous):

1| 1 -3 -6 8 | 1 -2 -8 ----------- 1 -2 -8 | 0 That's my synth. div. result

OpenStudy (anonymous):

when you look for zeros, it means to look for x's that would make this equation zero, which means solve the equation

OpenStudy (anonymous):

okay, yes see. i accidentally didn't computer my -6 and -2 right

OpenStudy (anonymous):

it's ok. then just solve \(x^2-2x-8=0\) as you would, as in factor it.

OpenStudy (anonymous):

so would my set up next be: x-1 1 -2 -8

OpenStudy (anonymous):

oh. okay.

OpenStudy (anonymous):

i think i need to do the quadratic formula since this isn't factoring right

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so for my answer I got x= +/- sqrt 19

OpenStudy (anonymous):

maybe i did something wrong, that isn't looking right either

OpenStudy (anonymous):

\(x^2-2x-8=(x-4)(x+2)\) ok?

OpenStudy (anonymous):

so i didn't need to do the formula?

OpenStudy (anonymous):

you need to factor \(x^2-2x-8\) do you know how to do that?

OpenStudy (anonymous):

so now that I have those zeros, 4 and -2 what do i do with them?

OpenStudy (anonymous):

for your (x-4)(x-2) when i just went to check it i got x^2-8x+8

OpenStudy (anonymous):

(x-4)(x\(+\)2) not MINUS

OpenStudy (anonymous):

x+2

OpenStudy (anonymous):

check again

OpenStudy (anonymous):

okay, sorry. yes i got it right this time

OpenStudy (anonymous):

so this is what you got \[x^4-3x^3-6x^2+8x=x(x-1)(x+2)(x-4)\]right?

OpenStudy (anonymous):

no, but that's oaky

OpenStudy (anonymous):

where did that x(x-1) pop up from?

OpenStudy (anonymous):

from the beginning...?

OpenStudy (anonymous):

oh, yeah i forgot that we took it out

OpenStudy (anonymous):

\[x^4-3x^3-6x^2+8x\]\[=x(x^3-3x^2-6x+8)\]\[=(x)(x-1)(x^2-2x-8)\]\[=(x)(x-1)(x-4)(x+2)\]ok?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then set \[x(x-1)(x+2)(x-4)=0\]and you should be able to find all of the solutions aka zeros.

OpenStudy (anonymous):

so my zeros would be 1, 4, -2 ?

OpenStudy (anonymous):

do you have the solutions?

OpenStudy (anonymous):

yes, x=0 x=1 x=-2 and x=4

OpenStudy (anonymous):

yes! YAY YOU GOT IT!

OpenStudy (anonymous):

do i need to do something with those? like put them in another synthetic division thing? when I plugged those answers into my computer they came back wrong

OpenStudy (anonymous):

no... your answers should be 0,1,-2,4...

OpenStudy (anonymous):

well thank you very much for your help sorry it took me so long to get it.

OpenStudy (anonymous):

it's ok and anytime!